OpenStudy (anonymous):

what is the solution to the rational equation (3/2)-(1/x-4)= -2/2x-8

OpenStudy (anonymous):

Is this written correctly? \[\frac{ 3 }{ 2 } - \frac{1}{x-4} = \frac{-2}{2x-8}\]

hero (hero):

I'm pretty sure I've seen this one before.

OpenStudy (anonymous):

Oh my D:

hero (hero):

Yup and I know what \[-\frac{2}{2x-8}\] reduces to

OpenStudy (anonymous):

To something...where...solutions are....

OpenStudy (anonymous):

...RATHER scarce indeed.